OpenStudy (anonymous):

Show all work in dividing y² – 3y + 2 over y² + 5y + 6 by y² + 2y – 8 over y² + y – 12. Use complete sentences to explain how to simplify this expression.

OpenStudy (mertsj):

\[\frac{y^2-3y+2}{y^2+5y+6}\div\frac{y^2+2y-8}{y^2+y-12}=\frac{(y-2)(y-1)}{(y+3)(y+2)}\times\frac{(y+4)(y-3)}{(y+4)(y-2)}\]

OpenStudy (mertsj):

\[\frac{(y-1)(y-3)}{(y+3)(y+2)}\]

OpenStudy (anonymous):

thanks!